# SeeGround: See and Ground for Zero-Shot Open-Vocabulary 3D Visual Grounding

### v1

@@ 3.4 Fusion Alignment Module
Although the 2D rendered images and text-based spatial descriptions provide substantial spatial information for SeeGround, directly inputting text and images without explicit processing can fail to associate 2D visual features with 3D spatial data. For instance, in scenes with multiple similar objects (e.g., several chairs), the model may struggle to link an object in the image with its corresponding description, leading to grounding errors. To address this, we introduce the Fusion Alignment Module, which explicitly associates key visual features in the scene with the textual description, ensuring a clear correspondence between the 2D rendered image and the text-based spatial descriptions.
Depth-Aware Visual Prompting. Specifically, after generating the rendered image $\mathcal{I}$, the OLT is used to retrieve the bounding box of each candidate object and extract the 3D points belonging to the object. These points are then projected onto the 2D image plane using the previously saved camera parameters $\mathcal{R}_{c}$ and $\mathcal{T}_{c}$, and visual markers are placed at the projection locations as prompts.
The simplest approach is to place these markers at the center of the projected points. However, due to occlusions caused by projecting 3D points onto a 2D plane, the projected center may correspond to other objects, misleading the model’s understanding of the scene. To address this issue, we leverage depth information to resolve occlusion problems. For each projected point, its depth is compared with the scene’s depth map to determine visibility. Only visible points are used to place the visual prompts. The visibility of an object $o$ is determined by checking whether enough of its constituent scene points are visible.
@@
After identifying the visible scene points, we place the visual prompts $\mathcal{M}_{o}$ at the projected center of these visible points, ensuring that the prompts reflect the unobstructed positions of the objects. The depth-aware visual prompting is expressed as follows:
$\mathcal{I}_{m}=\mathcal{I}\odot\left(1-\mathbf{1}_{\mathcal{P}_{\text{visible}}(o)}\right)+\mathcal{M}_{o}\odot\mathbf{1}_{\mathcal{P}_{\text{visible}}(o)},$ (6)
where $\mathbf{1}_{\mathcal{P}_{\text{visible}}(o)}$ is an indicator for the visibility of object $o$, and element-wise multiplication $\odot$ is used to apply these prompts selectively. The visualization of a typical example of $\mathcal{I}_{m}$ is shown in Fig. 4(b).
Object Prediction with 2D-VLM. Finally, given a query $\mathcal{Q}$, a rendered image $\mathcal{I}_{m}$, and the scene’s text description $\mathcal{T}$, the 2D-VLM predicts object $\hat{o}$ via:
$\hat{o}=\text{VLM}\left(\mathcal{Q}\,|\,\mathcal{I}_{m},\mathcal{T}\right).$ (7)
@@ 4.3 Ablation Study
Effect of View Selection Strategy. Tab. 4 demonstrates the advantage of our dynamic perspective alignment strategy (Fig. 3) over static ones. Static views – Center2Corner, Edge2Center, and Corner2Center – lack flexibility and struggle in complex scenes. Bird’s Eye View, though comprehensive, cannot adjust to the query and misses key spatial details like object orientation and height. By dynamically adjusting perspective based on the query, our method shows consistent improvement, particularly in “Hard” ($4.4\%$) and “Dependent” ($5.7\%$). This result underscores the importance of flexible and context-aware view selection strategy in 3D scene understanding.
@@
Robustness Evaluation with Incomplete Textual Descriptions. As shown in Fig. 7, we tested our approach’s robustness with incomplete textual information, simulating common misdetection scenarios. By omitting an anchor object from the text while retaining the target, our model uses visual cues to compensate, achieving accurate localization. In contrast, LLM performance degrades without the anchor. These results demonstrate that our method maintains high accuracy with partial text, underscoring the importance of integrating visual and textual data for reliable 3DVG.
@@
Figure 8: Error distributions between the Text-Only Method (a) and Our Method (b), based on four error types: Relation errors (Rel., spatial relationship misunderstandings like “next to” or “on the corner”), Classification errors (Cls., object category misidentifications), Viewpoint errors (View, errors in interpreting specific observation viewpoints), and Localization errors (Loc., errors in accurately pinpointing target object within the scene).
Type-Wise Error Analysis. To assess the potential limitations of our framework and guide future improvements, we conducted an error analysis on $185$ randomly selected samples across $10$ rooms, manually reviewing predictions to identify key error sources and guide improvements (see Fig. 8). The reduction in spatial localization and target identification errors highlights the importance of visual input in spatial understanding and object recognition. However, despite visual input, our method still has a high error rate in spatial relationships ($19\%$), indicating that precise spatial reasoning remains a challenge. Future work could benefit from advanced spatial reasoning modules. Current viewpoint selection strategies also fall short in handling complex scenarios like “when the window is on the left” or “upon entering from the door.” Finally, high-quality rendering provides clearer information about object boundaries, textures, and colors, helping models more accurately identify and distinguish objects, Our current use of point clouds from the dataset limits rendering precision.
@@ 5 Conclusion
Appendix
@@ 6.2 Details of Textual Prompt Design
These components are meticulously designed to guide the assistant in leveraging both visual and descriptive information for object identification. The structured format not only ensures clarity and consistency in the assistant’s responses but also facilitates effective communication and decision-making. By providing a comprehensive framework, the prompts enable the assistant to perform complex identification tasks with precision and reliability, which is critical in applications requiring high accuracy and interpretability.
In the main paper, we also discuss the process of determining anchor and target objects based on the query description. To further clarify this process, we provide an illustrative example of the prompt in Tab. 7. This prompt guides the VLM in identifying both the target object and its associated anchor object by analyzing their spatial and semantic relationships as described in the query. This design ensures the model focuses on identifying key objects while maintaining alignment with the query’s context.
@@ 6.3 Look-At-View Transform
In the Perspective Adaptation Module, we utilize the look_at_view_transform function to compute the extrinsic parameters of the virtual camera.
@@
    The look_at_view_transform function provides a systematic way to compute the extrinsic parameters of a camera in 3D space. The rotation matrix $\mathbf{R}_{c}$ transforms world coordinates into the camera’s view, while the translation vector $\mathbf{T}_{c}$ represents the camera’s position. These parameters are essential for rendering and aligning 3D scenes to match the desired perspective.
  • Query-Aligned Image Rendering. Once $\mathbf{R}_{c}$ and $\mathbf{T}_{c}$ are computed, the 3D scene $\mathbf{S}$ is projected into a 2D image plane to render the query-aligned image:
    $I=\texttt{Render}(\mathbf{S},\mathbf{R}_{c},\mathbf{T}_{c}).$ (13)
@@ 6.4 Details of Depth-Aware Visual Prompting
  • Generating Visual Prompts. To create visual prompts, we first retrieve the 3D bounding boxes and associated point sets for candidate objects from the OLT. These 3D points are then projected onto the 2D image plane using the camera parameters $\mathcal{R}_{c}$ (rotation matrix) and $\mathcal{T}_{c}$ (translation vector) obtained during the rendering process. Specifically, for a given 3D point $p=(x,y,z)$, the 2D projection $p^{\prime}=(x^{\prime},y^{\prime})$ is computed as:
    $p^{\prime}=\mathcal{R}_{c}p+\mathcal{T}_{c}.$ (14)
@@
    $\text{Visible}(o)=\begin{cases}1,&\text{if }\sum_{p_{o}\in\mathcal{P}_{o}}\text{Visible}(p_{o})\geq\alpha\cdot|\mathcal{P}_{o}|\\ 0,&\text{otherwise},\end{cases}$ (17)
@@ 7.2 Analysis of Visual Prompt Types
7.3 Real-world Image vs Rendered Image
@@
Figure 10: Illustration of SeeGround’s capability to resolve ambiguities in 3D visual grounding task. Incorrectly identified objects (Orange) and correctly identified objects (Green) are indicated to differentiate prediction accuracy, key cues are underlined.

### v2

@@ 3.3 Fusion Alignment Module
Although the 2D rendered images and text-based spatial descriptions provide substantial spatial information for SeeGround, directly inputting text and images without explicit processing can fail to associate 2D visual features with 3D spatial data. For instance, in scenes with multiple similar objects (e.g., several chairs), the model might struggle to link an object in the image with its corresponding description, leading to grounding errors. To address this, we introduce the Fusion Alignment Module, which explicitly associates key visual features in the scene with the textual description, ensuring a clear correspondence between the 2D rendered image and the text-based spatial descriptions.
Depth-Aware Visual Prompting. Specifically, after generating the rendered image $\mathbf{I}$, the object lookup table $\mathcal{OLT}$ retrieves the bounding box of each candidate object and extracts the 3D points belonging to the object. These points are then projected onto the 2D image plane using the precomputed camera parameters $\mathbf{R}_{c}$ and $\mathbf{T}_{c}$, and visual markers are placed at the projection locations as prompts.
The simplest approach is to place these markers at the center of the projected points. However, due to occlusions inherent in the 3D-to-2D projection process, the projected center may correspond to other objects, misleading the model’s understanding of the scene. To address this issue, we leverage depth information to resolve occlusion problems. For each projected point, its depth is compared with the scene’s depth map to determine visibility. Only visible points are used to place the visual prompts. The visibility of an object $o$ is determined by evaluating the proportion of its projected points that remain visible.
@@
After identifying the visible scene points, we place the visual prompt $\mathcal{M}_{o}$ of object $o$ at the projected center of these visible points, ensuring that the prompts reflect the unobstructed positions of the objects. The depth-aware visual prompting is expressed as follows:
$\mathbf{I}_{m}=\mathbf{I}\odot\left(1-\mathds{1}_{\mathcal{P}_{\mathrm{visible}}(o)}\right)+\mathcal{M}_{o}\odot\mathds{1}_{\mathcal{P}_{\mathrm{visible}}(o)},$ (6)
where $\mathds{1}_{\mathcal{P}_{\mathrm{visible}}(o)}$ is an indicator for the visibility of object $o$, and element-wise multiplication $\odot$ is used to apply these prompts selectively. The visualization of a typical example of $\mathbf{I}_{m}$ is shown in Fig. 4(b).
Object Prediction with 2D-VLM. Finally, given a query $\mathsf{Q}$, a rendered image $\mathbf{I}_{m}$, and the scene’s text description $\mathcal{T}$, the 2D-VLM predicts object $\hat{o}$ via:
$\hat{o}=\operatorname{VLM}\left(\mathsf{Q}\,|\,\mathbf{I}_{m},\mathcal{T}\right).\vspace{-0.2cm}$ (7)
@@ 4.3 Ablation Study
Effect of View Selection Strategy. Tab. 4 demonstrates the advantage of our dynamic perspective alignment strategy (Fig. 3) over static ones. Static views – Center2Corner, Edge2Center, and Corner2Center – lack flexibility and struggle in complex scenes. Bird’s Eye View, though comprehensive, cannot adjust to the query and misses key spatial details like object orientation and height. By dynamically adjusting perspective based on the query, our method shows consistent improvement, particularly in “Hard” ($4.4\%$) and “Dependent” ($5.7\%$). This result underscores the importance of a flexible and context-aware view selection strategy in 3D scene understanding.
@@
Robustness Evaluation with Incomplete Textual Descriptions. As shown in Fig. 7, we tested our approach’s robustness with incomplete textual information, simulating common misdetection scenarios. By omitting an anchor object from the text while retaining the target, our model uses visual cues to compensate, achieving accurate localization. In contrast, LLM performance degrades without the anchor. These results demonstrate that our method maintains high accuracy with partial text, underscoring the importance of integrating visual and textual data for more reliable 3DVG.
@@
Figure 8: Error distributions between the Text-Only Method (a) and Our Method (b), based on four error types: Relation Errors (Rel., spatial relationship misunderstandings like “next to” or “on the corner”), Classification Errors (Cls., object category misidentifications), Viewpoint Errors (View, errors in interpreting specific observation viewpoints), and Localization Errors (Loc., errors in pinpointing the target object within the scene).
Type-Wise Error Analysis. To assess the potential limitations of our framework and guide future improvements, we conducted an error analysis on $185$ randomly selected samples across $10$ rooms, manually reviewing predictions to identify key error sources and guide improvements (see Fig. 8). The reduction in spatial localization and target identification errors highlights the importance of visual input in spatial understanding and object recognition. However, despite visual input, our method still has a high error rate in spatial relationships ($19\%$), indicating that precise spatial reasoning remains a challenge. Future work could benefit from advanced spatial reasoning modules. Current viewpoint selection strategies also fall short in handling complex scenarios like “when the window is on the left” or “upon entering from the door”. Finally, high-quality rendering provides clearer information about object boundaries, textures, and colors, helping models more accurately identify and distinguish objects, Our current use of point clouds from the dataset limits rendering precision.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62306257), and the Guangzhou Municipal Science and Technology Project (No. 2024A03J0619 & 2024A04J4390). This work is also supported by the LASERi of HKUST(GZ), the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008) and Education Bureau of Guangzhou Municipality. This work is also supported by the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant No. M23L7b0021). Lingdong Kong is supported by the Apple Scholars in AI/ML Ph.D. Fellowship program.
The authors would like to sincerely thank the Program Chairs, Area Chairs, and Reviewers for the time and efforts devoted during the review process.
@@ 6.2 Details of Textual Prompt Design
These components are meticulously designed to guide the assistant in leveraging both visual and descriptive information for object identification. The structured format not only ensures clarity and consistency in the assistant’s responses but also facilitates effective communication and decision-making. By providing a comprehensive framework, the prompts enable the assistant to perform complex identification tasks with precision and reliability, which is critical in applications requiring high accuracy and interpret-ability.
In the main paper, we also discuss the process of determining anchor and target objects based on the query description. To further clarify this process, we provide an illustrative example of the prompt in Tab. 7. This prompt guides the VLM to identify both the target object and its associated anchor object by analyzing their spatial and semantic relationships as described in the query. This design ensures the model focuses on identifying key objects while maintaining alignment with the query’s context.
@@ 6.3 Look-At-View Transform
In the Perspective Adaptation Module, we utilize the $\operatorname{look\_at\_view\_transform}$ function to compute the extrinsic parameters of the virtual camera.
@@
    The $\operatorname{look\_at\_view\_transform}$ function provides a systematic way to compute the extrinsic parameters of a camera in 3D space. The rotation matrix $\mathbf{R}_{c}$ transforms world coordinates into the camera’s view, while the translation vector $\mathbf{T}_{c}$ represents the camera’s position. These parameters are essential for rendering and aligning 3D scenes to match the desired perspective.
  • Query-Aligned Image Rendering. Once $\mathbf{R}_{c}$ and $\mathbf{T}_{c}$ are computed, the 3D scene $\mathcal{S}$ is projected into a 2D image plane to render the query-aligned image $\mathbf{I}$:
    $\mathbf{I}=\operatorname{Render}(\mathcal{S},\mathbf{R}_{c},\mathbf{T}_{c}).$ (13)
@@ 6.4 Details of Depth-Aware Visual Prompting
  • Generating Visual Prompts. To create visual prompts, we first retrieve the 3D bounding boxes and associated point sets for candidate objects from the OLT. These 3D points are then projected onto the 2D image plane using the camera parameters $\mathbf{R}_{c}$ (rotation matrix) and $\mathbf{T}_{c}$ (translation vector) obtained during the rendering process. Specifically, for a given 3D point $p=(x,y,z)$, the 2D projection $p^{\prime}=(x^{\prime},y^{\prime})$ is computed as:
    $p^{\prime}=\mathbf{R}_{c}p+\mathbf{T}_{c}.$ (14)
@@
    $\text{Visible}(o)=\begin{cases}1,&\text{if }\sum_{p_{o}\in\mathcal{P}_{o}}\text{Visible}(p_{o})\geq\alpha\cdot|\mathcal{P}_{o}|,\\ 0,&\text{otherwise},\end{cases}$ (17)
@@ 7.2 Analysis of Visual Prompt Types
7.3 Results on Different Detectors.
Table 10: Performance comparison of different 3D detectors on the ScanRefer [5] validation set. Accuracy (Acc.) is reported for each method paired with different 3D detectors.
Method 3D Detector Acc.
ZSVG3D [66] Mask3D [49] $36.4$
OVIR-3D $19.3$
SeeGround Ground Truth $59.5$
Mask3D [46] $44.1$
OVIR-3D [38] $30.7$
Tab. 10 presents a performance comparison of different 3D detectors on the ScanRefer validation set, highlighting the impact of detector choice on grounding accuracy. With the same 3D detector (Mask3D), our method significantly outperforms the previous state-of-the-art approach, ZSVG3D, achieving an accuracy of $44.1$ compared to $36.4$. We also explore OVIR-3D as an alternative detector. The results show that our method achieves an accuracy of $30.7$ with OVIR-3D, while ZSVG3D achieves $19.3$ under the same setting. Additionally, the table reveals the upper-bound performance of our method when using ground truth (GT) proposals, reaching an accuracy of $59.5$. This underscores the importance of improving 3D object detection accuracy, as better detection directly translates to enhanced grounding results.
7.4 Real-world Image vs. Rendered Image
@@ 10 Public Resource Used
Figure 10: Illustration of SeeGround’s capability to resolve ambiguities in the 3D visual grounding task. Incorrectly identified objects (Orange) and correctly identified objects (Green) are indicated to differentiate prediction accuracy, key cues are underlined.